PREX Meeting Notes, Feb 25, 2010

Present: Bob Michaels, Ole Hansen, Mindy Jen, John LeRose, 
         Paul Souder, Krishna Kuamr, Dustin McNulty, Megan Friend
         Brian Quinn, Sasha Glamazdin, Alexandre Camsonne,
         Charles Hyde, Carlos Munoz Camacho, Vince Sulkosky

 GEMs - Ole

    A conference call was held yesterday.  Mechanical design
    and support to be fabricated in shops next week at UVa,
    Smith, and UMass.  Gas system being developed.  1 GEM to be
    setup in EEL for test.  List of electrical cables, connectors,
    and frontend boards: we have enough.  Software ~ready.


    Started installing ADCs for GEM on HRS.  Needs sequencers
    and some trigger items.  Will support setup in EEL
    Compton : will put EDET back in tomorrow.  Plane A repaired
    and shielding is optimized.  Want to restore vacuum and bake out,
    starting tomorrow.  Kink in photon line discussed: surveys
    being done, bigger pipe installed, some small movements in
    progress under direction of Sirish.

 Brian / CMU

     Question re: photon beamline kink and what was being done.

 Dustin McNulty

    - Aperture detectors for Moller checked out.  To be installed

    - Made 11 new cables for new dithering, have routed 
      them from BSY with Zafar's help.  
    - Will set up trigger.  S0 to be installed.  Alex and
      Samuel helping.

 Mindy Jen

     Presented electronics tests of CompSimple DAQ config
     (see HAPLOG).  A pulser is injected with typically 10
     coinc events per helicity period, together with a 
     random pulser at a few kHz.  Good signal/noise found.
     Alex asked:  What efficiency ?  (How many good coinc found ?)
     A suggestion was made to run the randoms up to very
     high rate and see if the DAQ breaks down.  (Of course
     that could happen with beam.)

 Sasha / Moller - 
     Vacuum leak neaer the target
     Additional NIM for aperture detector
     Working on fanning out signals to new and old DAQ
     Presented Moller issues to the Bteam meeting.

 Megan / Compton --

     Discussion of HAPLOG 2055 on improvements of
     analysis of HAPPEX-III data.

 Charles Hyde, Carlos Munoz Camacho, Alex Camsonne -- 
     Agreement to get the DVCS students to help with
     setup and checkout of base equipment, optics
     calibrations, Qsq, beam monitors, etc.  
     (good training for them)

 Discussion with Vince Sulkosky to help Dustin in as
     many ways as possible, e.g. with trigger setup.

 Next meeting:   Thurs Mar 4 at 3 pm

Index of Meetings